# MetaTrading: An Immersion-Aware Model Trading Framework for Vehicular Metaverse Services

### v1

@@ Abstract
Updates of extensive Internet of Things (IoT) data are critical to the immersion of vehicular metaverse services. However, providing high-quality and sustainable data in unstable and resource-constrained vehicular networks remains a significant challenge. To address this problem, we put forth a novel immersion-aware model trading framework that incentivizes metaverse users (MUs) to contribute learning models trained by their latest local data for augmented reality (AR) services in the vehicular metaverse, while preserving their privacy through federated learning. To comprehensively evaluate the contribution of locally trained learning models provided by MUs to AR services, we design a new immersion metric that captures service immersion by considering the freshness and accuracy of learning models, as well as the amount and potential value of raw data used for training. We model the trading interactions between metaverse service providers (MSPs) and MUs as an equilibrium problem with equilibrium constraints (EPEC) to analyze and balance their costs and gains. Moreover, considering dynamic network conditions and privacy concerns, we formulate the reward decisions of MSPs as a multi-agent Markov decision process. Then, a fully distributed dynamic reward method based on deep reinforcement learning is presented, which operates without any private information about MUs and other MSPs. Experimental results demonstrate that the proposed framework can effectively provide higher-value models for object detection and classification in AR services on real AR-related vehicle datasets compared to benchmark schemes.
@@ I Introduction
The metaverse, envisioned as one of the next evolution directions of the Internet, is designed to create a fully immersive, self-sustaining virtual environment in which people can engage in activities such as playing, working, and socializing [1]. This vision is propelled by advances in the fifth-generation (5G) and the forthcoming sixth-generation (6G) communication technologies, renowned for their low latency and high data throughput. These technologies play a critical role in seamlessly integrating the Internet of Things (IoT) data into metaverse services, thus bringing the once-fictional concept of immersive experiences closer to reality.
Metaverse services are beginning to reveal their vast potential across a broad spectrum of industries, from gaming and autonomous driving to education and marketing. Notably, the application of vehicles within the metaverse has captured significant interest due to the enhanced safety features and immersive experiences offered by state-of-the-art augmented reality (AR) technologies.
@@
Market report [2] forecasted that the global automotive metaverse market will grow from $1.9$ billion in 2022 to $6.5$ billion by 2030. Automakers like BMW have recently doubled down on AR technology in their projects. As shown in Fig. 2, augmented information makes driving safer by showing potential hazards hidden behind the vehicle in front of you. Moreover, Nissan’s [3] upcoming technology utilizes a 3D AR interface that merges the real world with the virtual world to provide the driver with augmented information about the surrounding area. In addition, Vanarama has also implemented a visual parking assistance function using AR [4], which is expected to identify empty spaces in busy car parks and display useful information, such as parking fees, on the windscreen.
@@
Motivations. An existing widely-used data collection method, as adopted by Nissan [6], involves collecting massive amounts of data through vehicle sensors, cameras, and roadside devices, and then processing all the data centrally. However, when it comes to the situation with multiple metaverse users (MUs) and metaverse service providers (MSPs) on behalf of different companies, the centralized data collection approach is not applicable and may lead to the following issues. First, AR services in the vehicular metaverse need to be highly immersive so that MUs feel fully immersed in the rendered environment, such as visualized driving. However, providing a seamless AR experience is challenging due to the latency caused by massive real-time data updates under unstable and resource-limited communication conditions. Note that the value of real-time data diminishes over time [7]. Also, delays can severely impact the MU’s experience and cause dizziness [8]. Second, the data to be transmitted may be sensitive and private, such as location, movement, and biometrics, which can create a better immersive experience but may inevitably increase the privacy risk of MUs [9]. Third, it is difficult to share data between MSPs due to competition. This leads to the same data being sampled multiple times by different MSPs in the physical world, resulting in wasted resources (e.g., sampling costs) and data redundancy. Moreover, tremendous data pressure and computational burden are imposed on MSPs. Therefore, an efficient method for collecting high-quality data while preserving privacy is needed to achieve immersive AR services.
Challenges. To address the above issues, we propose an immersion-aware model trading framework designed to provide data support for AR services and ensure privacy by integrating federated learning (FL). Adopting FL enables MUs to contribute diverse learning models to multiple MSPs using local data, thus effectively protecting sensitive information [10]. Moreover, learning models are uploaded to MSPs for AR services by individual MUs, rather than centrally transmitting large amounts of data, which significantly reduces the communication burden. Nonetheless, developing such a model trading framework still faces the following challenges:
  • MUs are typically self-interested and are reluctant to share learning models with MSPs due to the additional computation, communication, and energy overhead. Thus, the challenge here is how to incentivize them to become contributors in providing high-value learning models that will benefit vehicular metaverse services.
  • It is difficult to design an appropriate metric to comprehensively evaluate the value of learning models contributed by MUs due to the diverse aspects involved in model valuation.
  • MUs have different sampling costs and limited computation and communication resources, while MSPs differ in their model preferences and compete with each other. Hence, the next challenge is to model the dynamic competitive interactions among MSPs and to achieve an equilibrium of interests among MUs and MSPs in the model trading.
@@
Contributions. To address these challenges, we model the dynamic competitive trading interactions as an equilibrium problem with equilibrium constraints (EPEC) involving multiple leaders (MSPs) and multiple followers (MUs), which is a hierarchical optimization problem with equilibria at two levels [11]. Furthermore, to evaluate the value of learning models contributed by MUs to MSPs, we design a new metric called the immersion of the learning model (IoM). The metric jointly considers the freshness and accuracy of the learning model, as well as the amount and potential value of raw data used for training. Moreover, given the dynamic networks and the privacy concerns of MSPs, we formulate the reward decisions of MSPs as a multi-agent Markov decision process (MAMDP) and develop a multi-agent deep reinforcement learning (DRL)-based dynamic reward (MDDR) approach to obtain the reward decisions in a fully distributed manner. The outline of the proposed framework is shown in Fig. 2, and the main contributions of this paper are summarized as follows:
  • We propose an immersion-aware trading framework for learning models with FL assistance, incentivizing MUs to contribute high-value learning models for MSPs in a privacy-preserving and efficient manner. Moreover, a novel metric called “IoM” is designed to quantify the immersion enhancement provided by MUs for AR services.
  • We model the dynamic competitive trading interactions as an EPEC and provide theoretical support for the existence and uniqueness of equilibria at two levels. Moreover, a fully distributed MDDR approach is developed to adapt to dynamic environments and address complex reward decisions of MSPs without accessing any private information of MUs or MSPs.
  • We conduct extensive numerical simulations based on AR-related vehicle datasets to validate the efficacy and efficiency of MDDR and the proposed immersion-aware model trading framework. Compared to benchmark schemes, the framework better motivates MUs to provide higher-value learning models for object detection and classification functions of AR services effectively.
@@ II Related Work
Incentive Mechanism. Incentive Mechanisms have been widely applied to metaverse services [12, 13, 14, 15, 16, 17]. For example, Xu et al. [12] proposed an incentive framework for learning-based virtual reality (VR) services and designed a perceptual quality to measure VR users’ immersion in the virtual world. Sun et al. [13] investigated dynamic digital twins in air-assisted vehicular networking and designed a two-stage resource allocation incentive mechanism based on the Stackelberg game. Jiang et al. [14] developed a hierarchical game-theoretic coded distributed computing (CDC) framework for metaverse services. The framework employs coalition and Stackelberg games to select reliable CDC workers for MSPs and incentivize reliable CDC workers to participate in metaverse services. Xi et al. [16] proposed an incentive-based congestion control scheme for stochastic demand response in digital twin edge networks, using Lyapunov optimization and contract theory. Nguyen et al. [17] introduced MetaChain, an innovative blockchain-based framework designed to address the emerging challenges in developing metaverse applications. The framework integrates game theory to motivate MUs to contribute to the metaverse.
However, the above incentives for metaverse services mainly focus on digital twins, blockchain, or virtual reality, while ignoring the scenarios of multiple MSPs and multiple MUs in the vehicular metaverse. In this scenario, the complexity of interactions among MSPs and MUs requires novel incentive mechanisms to better facilitate the development and innovation of metaverse services.
AR and FL. The AR applications, characterized by smoothness and immersion requirements, provide immersive and interactive experiences for MUs by seamlessly integrating virtual content with the real world. For a better immersion experience, Cozzolino et al. [18] developed a latency-energy efficient task offloading for AR services and designed a best-effort resource provisioning algorithm to solve the problem of heterogeneous resource competition. Further, to reduce latency while protecting data privacy, Chen et al. [19] proposed a framework that integrates mobile edge computing and FL to solve the computational efficiency and low-latency object classification problem for AR services. In FL, MUs can train AI models on their local devices, thus enabling privacy-preserving collaborative machine learning while fully leveraging the computation resources of those MUs [20]. Furthermore, Lu et al. [21] investigated communication-efficient FL for digital twin edge networks in industrial IoT, which provides virtual and real synchronization while protecting privacy. Incorporating FL into AR services brings the benefits of reduced communication latency and privacy protection. However, the existing studies fail to consider the integration of FL into AR for the multi-MSP and multi-MU metaverse scenario, while overlooking the selfishness of MUs, i.e., whether they are willing to participate in FL learning.
Inspired by the above existing works, we design an immersion-aware framework integrated FL to provide a learning model trading platform for MSPs and MUs in the vehicular metaverse service markets, enabling MUs to contribute high-value learning models for AR services.
@@
Figure 3: The workflow of the immersion-aware model trading framework [^4] Image source: https://mixed-news.com/en/google-transforms-google-maps-into-the-backbone-of-our-ar-future/, https://about.facebook.com/meta, https://www.vanarama.com/blog/cars/4-ways-augmented-reality-will-revolutionise-the-automotive-industry..
@@ III System Model
We first present the system overview of the immersion-aware model trading framework in Section III-A. We next introduce the FL mechanism adopted by our framework in Section III-B and design the immersion metric of the learning model in Section III-C.
@@ III-A System Overview
The workflow of the immersion-aware model trading framework is illustrated in Fig. 4, aiming to provide MUs with immersive and interactive AR service experiences. MUs can contribute learning models to the AR services of MSPs for better immersive experiences. For example, services such as object detection can be improved by training on images of streets, pedestrians, and so on, captured by vehicles’ cameras. In this way, AR services can be better applied to visible driving, with more accurate and timely hazard warnings from windscreens. The framework mainly consists of MSPs owned by different companies (e.g., Meta, Nissan, and Google), an infrastructure layer, and an interaction layer with the MUs that enjoy vehicular metaverse services. MSPs are fueled by many technologies that work together to support vehicular metaverse services, including AI engines, AR/VR engines, metaverse trading engines, etc. The infrastructure layer with base stations (equipped with caches and edge servers) provides the basis of 5G/6G communication services for the interactions among MUs and MSPs. MUs play essential roles as both consumers and contributors to vehicular metaverse services.
TABLE I: Parameter Definitions and Notations
@@
$M$ Number of MUs
$N$ Number of MSPs
@@
$f_{mn}$ MU $m$’s computation resource used for MSP $n$
@@
The framework is a multi-engine configuration, and this paper focuses on two components: the trading engine and the AI engine of the MSP. The trading engine is responsible for incentivizing MUs to participate in mutually beneficial collaborations with MSPs, providing them with a trading platform to make their decisions (i.e., computation and communication resource allocation of MUs and reward decisions of MSPs). Then, each MU deciding to contribute utilizes the FL mechanisms integrated into AI engines to provide learning models for MSPs under the strategic guidance of the trading results. The specific workflow of the framework involves the following two phases.
Phase I (Incentive process): Initially, MSPs’ AI engines broadcast their global models to all MUs in the vehicular metaverse [^5] The FL global models owned by different MSPs we consider could be object detection or classification models for intelligent terrain mapping, intuitive road safety, visible driving, etc. Note that it is feasible to train on multiple tasks with the same local data. For example, for the same image, intuitive road safety focuses on pedestrians and vehicles, while intelligent terrain mapping focuses on landmarks and routes. (① in Fig. 4). Then, the trading engines of MSPs determine their digital currency prices per IoM, i.e., rewards to MUs, and broadcast them to MUs (②). Next, MUs determine their allocation of computation and communication resources for contributing learning models to various MSPs based on the IoM, costs, and rewards given by MSPs. (③). Based on the responses from MUs, the trading engines of MSPs adjust the digital currency prices to maximize their benefits. Steps ② and ③ are repeated until an agreement is reached among MUs and MSPs.
Phase II (FL process): Each MU, guided by decisions obtained through tradings, allocates computation resources to local training based on different global models from MSPs and generates local learning models for different MSPs (④). Then, MUs upload their updated learning models to the corresponding AI engines of MSPs by consuming communication resources (⑤). Following this, AI engines aggregate models from all MUs to update their global models and transmit them to their AR/VR engines for utilization (⑥). Finally, the AR/VR engines enable better AR services with the assistance of the AI engines (⑦). Steps ④-⑦ are repeated until the guidance time $T$ ends. To ease the presentation, we summarize some important notations in Table I.

### v2

@@ Abstract
Timely updating of Internet of Things (IoT) data is crucial for immersive vehicular metaverse services. However, challenges such as latency caused by massive data transmissions, privacy risks associated with user data, and computational burdens on metaverse service providers (MSPs) hinder continuous collection of high-quality data. To address these issues, we propose an immersion-aware model trading framework that facilitates data provision for services while ensuring privacy through federated learning (FL). Specifically, we first develop a novel multi-dimensional metric, the immersion of model (IoM), which assesses model value comprehensively by considering freshness and accuracy of learning models, as well as the amount and potential value of raw data used for training. Then, we design an incentive mechanism to incentivize metaverse users (MUs) to contribute high-value models under resource constraints. The trading interactions between MSPs and MUs are modeled as an equilibrium problem with equilibrium constraints (EPEC) to analyze and balance their costs and gains, where MSPs as leaders determine rewards, while MUs as followers optimize resource allocation. Furthermore, considering dynamic network conditions and privacy concerns, we formulate the reward decisions of MSPs as a multi-agent Markov decision process. To solve this, we develop a fully distributed dynamic reward algorithm based on deep reinforcement learning, without accessing any private information about MUs and other MSPs. Experimental results demonstrate that the proposed framework outperforms state-of-the-art benchmarks, achieving improvements in IoM of $38.3\%$ and $37.2\%$, and reductions in training time to reach the target accuracy of $43.5\%$ and $49.8\%$, on average, for the MNIST and GTSRB datasets, respectively.
@@ I Introduction
The metaverse, envisioned as a major evolutionary step for the Internet, aims to create a fully immersive, self-sustaining virtual environment for activities such as playing, working, and socializing [1]. This vision is propelled by advances in fifth- and sixth-generation (5G/6G) communication technologies, which offer low latency and high data throughput. These technologies play a critical role in seamlessly integrating the Internet of Things (IoT) data into metaverse services, thus bringing the once-fictional concept of immersive experiences closer to reality.
@@
Metaverse services are beginning to reveal their vast potential across a broad spectrum of industries, from gaming and autonomous driving to education and marketing. Notably, the application of vehicles within the metaverse [2, 3] has attracted significant interest, particularly for the enhanced traffic safety enabled by state-of-the-art augmented reality (AR) technologies. Market report [4] forecasted that the global automotive metaverse market will grow from $1.9$ billion in 2022 to $6.5$ billion by 2030. Automakers such as BMW have increasingly invested in AR technology. As shown in Fig. 2, augmented information makes driving safer by showing potential hazards hidden behind the vehicle in front of you. Moreover, Nissan’s [5] upcoming technology utilizes a 3D AR interface that merges the real world with the virtual world to provide the driver with augmented information about the surrounding area. In addition, Neuron mobility [^3] https://immersive-technology.com/augmentedreality/neuron-introduces-new-ar-parking-assistant/ has announced the launch of an innovative AR parking assistance system with the intention of improving passengers’ parking routines and trip-end experiences.
@@
Motivations. The widely used data collection method, as adopted by Nissan [7] and studied in [8], involves gathering massive amounts of data through vehicle sensors, cameras, and roadside devices, and all data being processed centrally. While this approach has been effective for various applications, when it comes to the situation with multiple metaverse users (MUs) and metaverse service providers (MSPs) associated with different companies, the centralized data collection approach is not applicable and may lead to the following issues. First, AR services in the vehicular metaverse need to be highly immersive so that MUs feel fully immersed in the rendered environment, such as visualized driving. However, the data synchronization will be hindered by the latency from massive real-time data updates under unstable and resource-limited communication conditions. Note that the value of real-time data diminishes over time [9]. Also, delays can severely impact the MU’s experience and cause dizziness [10]. Second, the data to be transmitted may be sensitive and private, such as location, movement, and biometrics, which can create a better immersive experience but may inevitably increase the privacy risk of MUs.
Federated learning (FL) has been adopted in prior work [11, 12, 6, 13] to enable collaborative model training without sharing raw data (e.g., sensor/imaging data from vehicles). In this approach, learning models are uploaded by individual MUs to MSPs for AR services, rather than transmitting large volumes of data centrally, thus significantly reducing the communication burden. However, MUs are typically self-interested and are reluctant to share learning models with MSPs due to the additional computation, communication, and energy overhead. To overcome this, incentive mechanisms using strategies such as contract theory [14], Stackelberg game [15, 16] and multi-winner sealed-bid auction [17], have been proposed to encourage MUs to contribute models. However, existing studies fail to explicitly assess the model value from multiple dimensions, making it difficult for both MSPs and MUs to quantify the benefits of the model for MSPs. Furthermore, most of these solutions are designed for a single MSP and do not consider the joint optimization of MUs’ limited computational and communication resources. As a result, an efficient and privacy-preserving framework for data synchronization is needed to improve the immersive AR experience in the vehicular metaverse.
To address the above research gaps, we propose an immersion-aware model trading framework for a multi-MSP, multi-MU vehicular metaverse. This framework is designed to incentivize MUs to become active contributors by providing models tailored to the specific needs of MSPs. However, MUs have different sampling costs and limited computational and communication resources, while MSPs differ in their model preferences and compete with each other. These inherent asymmetries and competitive dynamics make it challenging to establish an efficient model trading ecosystem. This leads to two key research questions that underpin our work:
  • How can MUs be effectively incentivized to contribute high-value learning models that benefit vehicular metaverse services for diverse MSPs?
  • How can the dynamic competitive interactions among MSPs be modeled to achieve a trade-off in gains between MUs and MSPs, ensuring sustainable and stable trading of learning models in the vehicular metaverse?
@@
Proposed Framework and Contributions. As depicted in Fig. 2, our model trading framework consists of four components: FL mechanism, metric design, game modeling, and dynamic algorithms. Specifically, we first design a new metric called the immersion of the learning model (IoM) to evaluate the value of learning models contributed by MUs to MSPs. This metric jointly considers the freshness and accuracy of the learning model, as well as the amount and potential value of raw training data. Building on this metric, we propose an immersion-aware incentive mechanism that aligns the interests of both MUs and MSPs. Then, we model the dynamic competitive trading interactions as an equilibrium problem with equilibrium constraints (EPEC), which is a hierarchical optimization problem with equilibria at two levels [18]. Moreover, given the dynamic networks and the privacy concerns of MSPs, we formulate the reward decisions of MSPs as a multi-agent Markov decision process (MAMDP) and develop a multi-agent deep reinforcement learning (DRL)-based dynamic reward (MDDR) approach to obtain the reward decisions in a fully distributed manner. In summary, the key contributions of this work are as follows.
  • Incentive mechanism design for immersion-aware model trading. From the perspectives of both MUs and MSPs, we propose an incentive mechanism that encourages MUs to contribute high-value learning models tailored to the specific demands of MSPs. To our knowledge, this is the first study focusing on incentive mechanism design for efficient and privacy-preserving data synchronization in multi-MU, multi-MSP vehicular metaverse environments.
  • Novel design of multi-dimensional metric. To quantify immersion enhancement provided by MUs for AR services, we design an immersion metric of the learning model integrating four critical dimensions: the freshness and accuracy of the learning model, as well as the amount and potential value of raw training data. Freshness is captured through age of information (AoI), while potential value is evaluated by the difference between model predictions and true labels. This metric enables fine-grained evaluation of model value under resource constraints and provides a basis for incentive and decision-making strategies in model trading.
  • Theoretical analysis and algorithm designs. Given multiple resource-constrained MUs and competing MSPs, we model their interactions as an equilibrium problem with equilibrium constraints (EPEC). We theoretically prove the existence and uniqueness of the equilibrium, and further develop a fully distributed MDDR approach that adapts to dynamic environments and operates without accessing any private information of MUs or MSPs.
  • Performance evaluation. We conduct extensive numerical simulations based on AR-related vehicle datasets to validate the efficacy and efficiency of MDDR and the proposed immersion-aware model trading framework. Numerical results show that our proposed mechanism improves the IoM by $38.3\%$ and $37.2\%$, and reduces the training time to reach the target accuracy by $43.5\%$ and $49.8\%$, on average, for the MNIST and GTSRB datasets, respectively, compared with state-of-the-art benchmarks.
@@ II Related Work
In this section, we survey work related to our study in terms of edge-enabled vehicular metaverse services, FL for AR, and incentive mechanisms for data synchronization in the metaverse.
II-A Edge-Enabled Vehicular Metaverse Services
Vehicular metaverse services require ultra-low latency and high reliability to enable immersive experiences such as AR navigation and real-time digital twins. To meet these demands, researchers have extensively explored edge-based strategies for task offloading and resource management. For instance, Feng et al. [19] proposed a resource allocation framework for AR-powered vehicular edge metaverses that aims to maximize overall system utility. Similarly, Khan et al. [20] introduced a cooperative framework integrating task offloading, sensing, learning, and communication to reduce transmission energy consumption and latency in resource-constrained environments. Beyond static offloading, Chen et al. [21] developed a multi-agent deep reinforcement learning approach for dynamic avatar task migration. In terms of security and privacy, Kang et al. [22] presented a cross-metaverse empowered dual pseudonym management framework to mitigate privacy leakage risks during the dynamic communications among vehicular edge metaverses. Moreover, an optimization framework for AR-assisted driving was proposed in [23] to enhance the robustness of metaverse maps against adversarial attacks. This framework maximizes the mean average precision (mAP) of adversarial patch detection while minimizing AR scene uplink latency and worst-case battery consumption. Recent studies have also explored applying FL to vehicular metaverse services as a means of preserving user privacy during distributed model training. These efforts, particularly in AR-enhanced metaverse environments, are discussed in the next section. To ensure service continuity and prevent user disengagement during unexpected failures, Qiu et al. [24] proposed deploying redundant backup virtual vehicle services and keeping the age of backup information to improve the reliability of services.
These works, from diverse perspectives, have significantly advanced the security, reliability, and performance of vehicular metaverse services, providing valuable insights for future research. Notably, distinct from existing studies, our work focuses on designing an incentive mechanism that transforms MUs into active contributors. These contributors efficiently provide high-value models to MSPs, thereby enhancing the overall service experience.
II-B Federated Learning for AR in the Metaverse
Privacy preservation is a critical concern in vehicular metaverse services, particularly when training models using sensitive user data. Federated learning (FL) has emerged as a promising technique to enable collaborative model training without sharing raw data. For example, Chen et al. [11] proposed a framework that integrates mobile edge computing and FL to solve the computational efficiency and low-latency object classification problem for AR services. However, they did not address how communication and computational resources should be allocated within the FL framework. To overcome this limitation, Zhou et al. [12] incorporated FL into mobile AR systems in the metaverse to enable collaborative model training. Their approach considers the trade-offs among energy consumption, execution latency, and model accuracy to meet varying demands across different application scenarios. To further improve learning efficiency, an adaptive and resource-efficient FL algorithm tailored for AR applications was proposed in [6]. This approach mitigates the effects of non-IID data distributions, reduces resource usage, and enhances the quality of experience. Additionally, Hazarika et al. [13] explored the integration of quantum computing with FL to provide a cost-efficient and adaptive solution for the dynamic nature of vehicular environments.
Incorporating FL into AR services brings the benefits of reduced communication latency and privacy protection. However, the existing studies fail to consider the integration of FL into AR for the multi-MSP and multi-MU metaverse scenario, while overlooking the selfishness of MUs, i.e., whether they are willing to participate in FL learning.
II-C Incentive Mechanism for Data Synchronization in the Metaverse.
The quality of metaverse services depends heavily on massive sensing data from the real world. To facilitate this, incentive mechanisms are widely employed to motivate MUs to contribute data and participate in model construction, thereby enhancing their service experience. For instance, Zhang et al. [25] proposed a vehicle-assisted data sensing framework that incentivizes vehicles to contribute sensing data to metaverse service, improving the driving experience for vehicle users. Similarly, Lin et al. [26] introduced a contract-based incentive design aimed at optimizing the long-term profits of digital twin service providers, addressing service congestion caused by stochastic demand responses. Xu et al. [27] developed a metaverse-based unmanned aerial vehicle (UAV) swarm system that uses digital twins and semantic communication to enhance data synchronization and reduce communication latency. Moreover, Lotfi et al. [28] introduced VMGuard, a novel four-layer security framework for the vehicular metaverse that protects against data poisoning attacks by using a reputation-based incentive mechanism. The framework leveraged user feedback and subjective logic modeling to assess the trustworthiness of devices. However, while these works successfully incentivize MUs to contribute raw data, they also pose a risk to user privacy due to potential data leakage.
To address the privacy concern, Kang et al. [14] developed a cross-chain FL framework with an AoI-based contract theory model under prospect theory to incentivize data sharing. They utilized a hierarchical cross-chain architecture with a main chain and multiple subchains to perform decentralized, privacy-preserving, and secure data training in both virtual and physical spaces. Baccour et al. [15] proposed a dual game-theoretic framework for federated meta-learning (FML) in metaverse services. They introduced a reputation system and a Stackelberg game-based incentive mechanism to enhance data privacy, minimize energy costs, and improve FML efficacy, outperforming traditional clustering methods in training performance and metaverse utility. Li et al. [16] designed a satisfaction metric that considers data size, AoI, and training latency. They integrated this metric into utility functions to incentivize node participation in FL while optimizing resource allocation. However, these works are designed for a single MSP scenario, limiting scalability and applicability in multi-server scenarios commonly encountered in large-scale metaverse environments.
Building on the concept of multiple MSPs, Zhang et al. [17] proposed a privacy-preserving auction mechanism for learning model as an NFT in Blockchain-driven Metaverse. They employed an imperfect information Stackelberg game to optimize the strategies of MUs and MSPs, and introduced a multi-winner sealed-bid auction mechanism to enhance the quality of FL-NFT. However, this approach primarily focuses on modeling NFTs and overlooks the joint optimization of computational and communication resources. Moreover, it fails to evaluate the model’s value from multiple dimensions.
Inspired by the above existing works, this study focuses on multi-MSP and multi-MU scenarios in vehicular metaverse services. By assigning MUs the dual roles of service consumers and model contributors, we build a low-cost, privacy-preserving data support framework. This allows MUs to enjoy immersive in-vehicle services (e.g., AR navigation, digital twin interactions) while participating as distributed model contributors in service optimization. This framework continuously incentivizes MUs to contribute high-value learning models for metaverse services, thereby enhancing the overall service experience.
@@ II-B Federated Learning for AR in the Metaverse
Figure 3: Workflow of the immersion-aware model trading framework [^5] Image source: https://mixed-news.com/en/google-transforms-google-maps-into-the-backbone-of-our-ar-future/, https://about.facebook.com/meta, https://www.vanarama.com/blog/cars/4-ways-augmented-reality-will-revolutionise-the-automotive-industry..
@@ III System Model
We first present the system overview of the immersion-aware model trading framework in Section III-A. Next, we introduce the FL mechanism adopted by our framework in Section III-B and design the immersion metric of the learning model in Section III-C.
@@ III-A System Overview
Fig. 5 illustrates the workflow of our immersion-aware model trading framework, designed to enhance the immersive and interactive AR experiences of MUs. MUs can contribute learning models to the AR services of MSPs for better immersive experiences. For example, services such as object detection can be improved by training on images of streets, pedestrians, and so on, captured by vehicles’ cameras. In this way, AR services can be better applied to visible driving, with more accurate and timely hazard warnings from windscreens. The framework mainly consists of MSPs owned by different companies (e.g., Meta, Nissan, and Google), an infrastructure layer, and an interaction layer with the MUs that enjoy vehicular metaverse services. MSPs are powered by various technologies supporting vehicular metaverse services, such as AI engines, AR/VR engines, metaverse trading engines. The infrastructure layer with base stations (equipped with caches and edge servers) provides the basis of 5G/6G communication services for the interactions among MUs and MSPs. MUs play essential roles as both consumers and contributors to vehicular metaverse services.
TABLE I: Definitions of Notation
@@
$m,M$ Index of an MU, Number of MUs
$n,N$ Index of an MSP, Number of MSPs
@@
$f_{mn}$ MU $m$’s computational resource used for MSP $n$
@@
The framework is a multi-engine configuration, and this paper focuses on two components: the trading engine and the AI engine of the MSP. The trading engine is responsible for incentivizing MUs to participate in mutually beneficial collaborations with MSPs, providing them with a trading platform to make their decisions (i.e., computational and communication resource allocation of MUs and reward decisions of MSPs). Then, each MU deciding to contribute utilizes the FL mechanisms integrated into AI engines to provide learning models for MSPs under the strategic guidance of the trading results. The specific workflow of the framework involves the following two phases.
Phase I (Incentive process): Initially, MSPs’ AI engines broadcast their global models to all MUs in the vehicular metaverse [^6] The FL global models owned by different MSPs we consider could be object detection or classification models for intelligent terrain mapping, intuitive road safety, visible driving, etc. Note that it is feasible to train on multiple tasks with the same local data. For instance, for the same image, intuitive road safety focuses on pedestrians and vehicles, while intelligent terrain mapping focuses on landmarks and routes. (① in Fig. 5). Then, the trading engines of MSPs determine their digital currency prices per IoM, i.e., rewards to MUs, and broadcast them to MUs (②). Next, MUs determine their allocation of computational and communication resources for contributing learning models to various MSPs based on the IoM, costs, and rewards given by MSPs (③). Based on the responses from MUs, the trading engines of MSPs adjust the digital currency prices to maximize their own utilities. Steps ② and ③ are repeated until an agreement is reached among MUs and MSPs.
Phase II (FL process): Each MU, guided by decisions obtained through tradings, allocates computational resources to local training based on different global models from MSPs and generates local learning models for different MSPs (④). Then, MUs upload their updated learning models to the corresponding AI engines of MSPs by consuming communication resources (⑤). Following this, AI engines aggregate models from all MUs to update their global models and transmit them to their AR/VR engines for utilization (⑥). Finally, the AR/VR engines provide enhanced AR services to MUs (⑦). Steps ④-⑦ are repeated until the guidance time $T$ ends. To ease the presentation, we summarize some important notations in Table I.